The revenues of a company increased by 31% in one year and decreased by 27% in year two. What is the overall change over the two-year period?


1
Expert's answer
2021-09-13T17:24:19-0400

Lets assume the Revenue of the company at time zero is 100

Then, its revenue at the end of Year 1 "= 100\times (1+0.31) =131"

Then, its revenue at the end of Year 2 "= 131\times(1-0.27) =131\times 0.73 = 95.63"

The Overall change in revenue over two year period "= 100-95.63 = 4.37\%" ( decrease in revenue)